# Gated Integration of Low-Rank Adaptation for Continual Learning of Language Models

### v1

@@ A.1 Gradient Projection Memory
We initialize the first $L$ layers of $g_{t}(\cdot)$ using the corresponding layers from the previous gating module $g_{t-1}(\cdot)$. Therefore, the first $L$ layers of $g_{t}(\cdot)$ can be viewed as being initialized at the beginning of the first task and continue their training until the $t$-th task. Additionally, the first $L$ layers in $g_{i}(\cdot)$ serve as checkpoints, preserving the state of $g_{t}(\cdot)$ after learning the $i$-th task ($1\leq i\leq t$). At this time, existing method gradient projection memory (GPM) DBLP:conf/iclr/SahaG021 can be used to learn matrices $\{\bm{M}_{t,l}\}_{l=1}^{L+1}$, where the columns of $\bm{M}_{t,l}$ approximate the orthonormal bases of the subspace $\mathcal{M}_{t,l}$. Specifically, when $t=1$, since there is no old task, ${\mathcal{M}}_{1,l}$ is a null space and $\bm{M}_{1,l}$ is a zero matrix. After learning the $t$-th new task, GPM expands ${\mathcal{M}}_{t,l}$ to ${\mathcal{M}}_{t+1,l}$ by first computing the input matrix $\bm{H}_{t,l}$ where each column of $\bm{H}_{t,l}$ represents an input to the $l$-th layer. Then, the component of $\bm{H}_{t,l}$ already in $\mathcal{M}_{t,l}$ is removed by
@@ A.2 More Details of O-LoRA and InfLoRA
O-LoRA wang2023orthogonal ensures that the new LoRA branch remains orthogonal to all the old LoRA branches. Specifically, during the learning of the $t$-th new task with the $t$-th LoRA branch ($\bm{A}_{t},\bm{B}_{t}$), O-LoRA computes the inner product between the new and old LoRA branches as
@@
$\frac{1}{|\mathcal{D}_{t}|}\sum_{(\bm{x}_{t},\bm{y}_{t})\in\mathcal{D}_{t}}\sum_{j=1}^{|\bm{y}_{t}|}\log\left[P(y_{t,j}|\bm{x}_{t},y_{t,1},...,y_{t,j-1})\right]+\lambda\sum_{i=1}^{t-1}\sum_{j,k}||\bm{O}_{i,t}[j,k]||_{2}^{2}$ (15)
For further details on O-LoRA, we refer readers to the original paper wang2023orthogonal .
@@
InfLoRA liang2024inflora ensures orthogonality between the new LoRA branch and the gradients of old tasks. Specifically, it shows that only fine-tuning the down-projection matrix $\bm{A}_{t}$ in the new LoRA branch is equivalent to directly fine-tuning the pre-trained weights within a subspace spanned by the rows of $\bm{B}_{t}$. Therefore, before learning the $t$-th task, InfLoRA designs $\bm{B}_{t}$ to be orthogonal to the gradients of the old tasks. During the learning of the $t$-th task, InfLoRA only tunes $\bm{A}_{t}$ in the new LoRA branch while freezing $\bm{B}_{t}$ and all the old LoRA branches. For further details on InfLoRA, we refer readers to the original paper liang2024inflora .
@@
If the constraints in (6) are satisfied, subspaces $\{\mathcal{M}_{t,l}\}_{l=1}^{L+1}$ remain unchanged during the learning of the $t$-th task. Furthermore, for any input $\bm{x}$ from the previous $t-1$ tasks, $g_{t}(\bm{x})$ remains unchanged during the learning of the $t$-th task.
@@
For any $\bm{x}$ from previous $t-1$ tasks, we rewrite $\bm{g}_{t}(\bm{x})$ as
$g_{t}(\bm{x})=f(\bm{G}_{t,L+1}\bm{p}_{L}),$
$\bm{p}_{l}=\sigma(\bm{G}_{t,l}\bm{p}_{l-1}),~{}l\in\{1,2,...,L\},$
$\bm{p}_{0}={\rm Pool}({\rm Token}(\bm{x})).$ (16)
Since $\bm{p}_{0}={\rm Pool}({\rm Token}(\bm{x}))$ is unrelated to the parameters of the new gating module $g_{t}(\cdot)$, $\bm{p}_{0}$ does not change with the update of $g_{t}(\cdot)$. Since $\mathcal{M}_{t,1}$ is spanned by $\bm{p}_{0}$, $\mathcal{M}_{t,1}$ remains unchanged during the learning of the $t$-th task.
Suppose that we have proven that $\bm{p}_{l-1}$ does not change with the update of the new gating module $g_{t}(\cdot)$ ($1\leq l\leq L$). Since $\mathcal{M}_{t,l}$ is spanned by $\bm{p}_{l-1}$, $\mathcal{M}_{t,l}$ remains unchanged during the learning of the $t$-th task. At this point, $\bm{p}_{l}$ can be expressed as
$\bm{p}_{l}=\sigma(({\rm Init}(\bm{G}_{t,l})+\Delta\bm{G}_{t,l})\bm{p}_{l-1})=\sigma({\rm Init}(\bm{G}_{t,l})\bm{p}_{l-1}).$ (17)
Here, the second equality holds since $\bm{p}_{l-1}\in\mathcal{M}_{t,l}$ and $\Delta\bm{G}_{t,l}\bot\mathcal{M}_{t,l}$. Therefore, $\bm{p}_{l}$ does not change with the update of the new gating module $g_{t}(\cdot)$ ($1\leq l\leq L$). Since $\mathcal{M}_{t,l+1}$ is spanned by $\bm{p}_{l}$, $\mathcal{M}_{t,l+1}$ remains unchanged during the learning of the $t$-th task.
@@
$\bm{g}_{t}(\bm{x})=f(({\rm Init}(\bm{G}_{t,L+1})+\Delta\bm{G}_{t,L+1})\bm{p}_{L})=f({\rm Init}(\bm{G}_{t,L+1})\bm{p}_{L}).$ (18)
Here, the second equality holds since $\bm{p}_{L}\in\mathcal{M}_{t,L+1}$ and $\Delta\bm{G}_{t,L+1}\bot\mathcal{M}_{t,L+1}$. ∎
@@ Appendix B More Details of Experimental Settings
B.1 Evaluation Metrics
We use ${\rm A}_{j,i}$ to denote the model’s performance on the $i$-th task once the model learns the $j$-th task. Specifically, ${\rm A}_{j,i}$ represents accuracy for classification tasks and Rouge-L lin2004rouge for other types of tasks. Following traditional CL works chaudhry2020continual ; chaudhry2019tiny , we employ average performance (AP) and forgetting (FT) to evaluate the model’s performance. The formulas for these two metrics are defined as
B.2 More Details of Datasets
@@
Task639__\__multi__\__woz__\__user__\__utterance__\__generation summarization Rouge-L
Task1590__\__diplomacy__\__text__\__generation summarization Rouge-L
Task1729__\__personachat__\__generate__\__next summarization Rouge-L
Task181__\__outcome__\__extraction information extraction Rouge-L
Task748__\__glucose__\__reverse__\__cause__\__event__\__detection information extraction Rouge-L
Task1510__\__evalution__\__relation__\__extraction information extraction Rouge-L
Task002__\__quoref__\__answer__\__generation dialogue generation Rouge-L
Task073__\__commonsenseqa__\__answer__\__generation dialogue generation Rouge-L
Task591__\__sciq__\__answer__\__generation dialogue generation Rouge-L
Task511__\__reddit__\__tifu__\__long__\__text__\__summarization question answering Rouge-L
Task1290__\__xsum__\__summarization question answering Rouge-L
Task1572__\__samsum__\__summary question answering Rouge-L
Task363__\__sst2__\__polarity__\__classification sentiment analysis Accuracy
Task875__\__emotion__\__classification sentiment analysis Accuracy
Task1687__\__sentiment140__\__classification sentiment analysis Accuracy
@@
SuperNI Benchmark 1 task1572 $\rightarrow$ task363 $\rightarrow$ task1290 $\rightarrow$ task181 $\rightarrow$ task002 $\rightarrow$ task1510 $\rightarrow$ task639 $\rightarrow$ task1729 $\rightarrow$ task073 $\rightarrow$ task1590 $\rightarrow$ task748 $\rightarrow$ task511 $\rightarrow$ task591 $\rightarrow$ task1687 $\rightarrow$ task875
2 task748 $\rightarrow$ task073 $\rightarrow$ task1590 $\rightarrow$ task639 $\rightarrow$ task1572 $\rightarrow$ task1687 $\rightarrow$ task591 $\rightarrow$ task363 $\rightarrow$ task1510 $\rightarrow$ task1729 $\rightarrow$ task181 $\rightarrow$ task511 $\rightarrow$ task002 $\rightarrow$ task1290 $\rightarrow$ task875
CL Benchmark 3 MNLI $\rightarrow$ CB $\rightarrow$ WiC $\rightarrow$ COPA $\rightarrow$ QQP $\rightarrow$ BoolQA $\rightarrow$ RTE $\rightarrow$ IMDB $\rightarrow$ Yelp $\rightarrow$ Amazon $\rightarrow$ SST-2 $\rightarrow$ DBpedia $\rightarrow$ AG News $\rightarrow$ MultiRC $\rightarrow$ Yahoo
4 Yelp $\rightarrow$ Amazon $\rightarrow$ MNLI $\rightarrow$ CB $\rightarrow$ COPA $\rightarrow$ QQP $\rightarrow$ RTE $\rightarrow$ IMDB $\rightarrow$ SST-2 $\rightarrow$ DBpedia $\rightarrow$ AG News $\rightarrow$ Yahoo $\rightarrow$ MultiRC $\rightarrow$ BoolQA $\rightarrow$ WiC
B.3 More Implementation Details
Following existing CL works ouyang2022training ; wang2023orthogonal ; wei2021finetuned , all methods are implemented using instruction tuning ouyang2022training . Experiments are conducted on 5 NVIDIA RTX A6000 GPUs with AdamW DBLP:conf/iclr/LoshchilovH19 as the optimizer. The type of CPU is Intel(R) Xeon(R) Gold 6240R CPU @ 2.40GHz. For T5-Large and T5-XL, their relatively smaller model sizes allow experiments to be performed on a single A6000 GPU with gradient accumulation. For Llama-2-7B and Llama-2-13B, data parallelism with DeepSpeed ZeRO-2 rasley2020deepspeed is prioritized across multiple A6000 GPUs. FlashAttention-2 daoflashattention is employed to reduce memory usage during training, ensuring sufficient GPU memory to enable DeepSpeed ZeRO-2 whenever possible. However, if the sequence lengths of certain tasks are too long to enable DeepSpeed ZeRO-2 even with FlashAttention-2, DeepSpeed ZeRO-3 is utilized to handle these tasks.
To ensure fair comparisons, for all the methods based on LoRA, we follow existing CL methods hu2021lora ; wang2023orthogonal ; zhao2024sapt by integrating the LoRA architecture into the query and value components of the multi-head attention mechanism in each Transformer block. Following existing works wang2023orthogonal ; zhao2024sapt , for all the methods based on LoRA, the rank of a single LoRA branch is set to 4 for Order 1 and Order 2, and 8 for Order 3 and Order 4. We also vary the rank in LoRA branches and show the results in Appendix C.5.
@@ B.3 More Implementation Details
B.4 More Details about the Architecture of the Gating Module
@@
$g_{i}(\bm{x})=f(\bm{G}_{i,L+1}\bm{p}_{L}),$
$\bm{p}_{l}=\sigma(\bm{G}_{i,l}\bm{p}_{l-1}),~{}l\in\{1,2,...,L\},$
$\bm{p}_{0}={\rm Pool}({\rm Token}(\bm{x})).$ (20)
Non-linear activation function $\sigma(\cdot)$ is set to SiLU elfwing2018sigmoid . For all experiments, unless otherwise stated, $L$ is set to 2. In other words, the gating module $g_{i}(\cdot)$ has three layers. For T5-Large and T5-XL, the parameters in the $i$-th gating module $g_{i}(\cdot)$ are $\bm{G}_{i,1}\in\mathbb{R}^{100\times d}$, $\bm{G}_{i,2}\in\mathbb{R}^{d\times 100}$ and $\bm{G}_{i,3}\in\mathbb{R}^{1\times d}$. For Llama-2-7B and Llama-2-13B, the parameters in the $i$-th gating module $g_{i}(\cdot)$ are $\bm{G}_{i,1}\in\mathbb{R}^{50\times d}$, $\bm{G}_{i,2}\in\mathbb{R}^{d\times 50}$ and $\bm{G}_{i,3}\in\mathbb{R}^{1\times d}$. Here, $d$ denotes the dimension of the embeddings. For different models, $d$ is 1024 for T5-Large and T5-XL, 4096 for Llama-2-7B, and 5120 for Llama-2-13B.
@@
B.5 Computation of Trainable Parameters
@@
B.5.1 Computation of Trainable Parameters in T5-Large
In T5-Large, the projection weights for the query and value components have shapes $\bm{W}_{q},\bm{W}_{v}\in\mathbb{R}^{1024\times 1024}$. The model consists of 24 self-attention modules in the encoder, 24 self-attention modules in the decoder, and 24 cross-attention modules in the decoder, resulting in a total of $(24+24+24)*2=144$ pre-trained weights that incorporate the LoRA architecture.
During the learning of the $t$-th new task, O-LoRA updates the parameters $\bm{A}_{t}\in\mathbb{R}^{1024\times r}$ and $\bm{B}_{t}\in\mathbb{R}^{r\times 1024}$, resulting in $1024*r*144+r*1024*144=294912r$ trainable parameters. When $r=4$, the number of trainable parameters in O-LoRA is $294912*4=1179648=1.18$M. InfLoRA only updates the parameters $\bm{A}_{t}\in\mathbb{R}^{1024\times r}$, resulting in $1024*r*144=147456r$ trainable parameters. When $r=4$, the number of trainable parameters in InfLoRA is $0.59147456r=589824=$ 0.59M.
GainLoRA introduces an additional new gating module $g_{t}(\cdot)$ with parameters $\bm{G}_{t,1}\in\mathbb{R}^{100\times 1024}$, $\bm{G}_{t,2}\in\mathbb{R}^{1024\times 100}$ and $\bm{G}_{t,3}\in\mathbb{R}^{1\times 1024}$. Therefore, the number of trainable parameters in GainLoRA (O-LoRA) is $1179648+1024*100+1024*100+1024=1385472=1.39$M. The number of trainable parameters in GainLoRA (InfLoRA) is $589824+1024*100+1024*100+1024=795648=0.80$M.
B.5.2 Computation of Trainable Parameters in T5-XL

### v2

@@ A.1 Gradient Projection Memory
We initialize the first $L$ layers of $g_{t}(\cdot)$ using the corresponding layers from the previous gating module $g_{t-1}(\cdot)$. Therefore, the first $L$ layers of $g_{t}(\cdot)$ can be viewed as being initialized at the beginning of the first task and continue their training until the $t$-th task. Additionally, the first $L$ layers in $g_{i}(\cdot)$ serve as checkpoints, preserving the state of $g_{t}(\cdot)$ after learning the $i$-th task ($1\leq i\leq t$). At this time, existing method gradient projection memory (GPM) [47] can be used to learn matrices $\{\bm{M}_{t,l}\}_{l=1}^{L+1}$, where the columns of $\bm{M}_{t,l}$ approximate the orthonormal bases of the subspace $\mathcal{M}_{t,l}$. Specifically, when $t=1$, since there is no old task, ${\mathcal{M}}_{1,l}$ is a null space and $\bm{M}_{1,l}$ is a zero matrix. After learning the $t$-th new task, GPM expands ${\mathcal{M}}_{t,l}$ to ${\mathcal{M}}_{t+1,l}$ by first computing the input matrix $\bm{H}_{t,l}$ where each column of $\bm{H}_{t,l}$ represents an input to the $l$-th layer. Then, the component of $\bm{H}_{t,l}$ already in $\mathcal{M}_{t,l}$ is removed by
@@ A.2 More Details of O-LoRA and InfLoRA
O-LoRA [61] ensures that the new LoRA branch remains orthogonal to all the old LoRA branches. Specifically, during the learning of the $t$-th new task with the $t$-th LoRA branch ($\bm{A}_{t},\bm{B}_{t}$), O-LoRA computes the inner product between the new and old LoRA branches as
@@
[…]
For further details on O-LoRA, we refer readers to the original paper [61].
@@
InfLoRA [32] ensures orthogonality between the new LoRA branch and the gradients of old tasks. Specifically, it shows that only fine-tuning the down-projection matrix $\bm{A}_{t}$ in the new LoRA branch is equivalent to directly fine-tuning the pre-trained weights within a subspace spanned by the rows of $\bm{B}_{t}$. Therefore, before learning the $t$-th task, InfLoRA designs $\bm{B}_{t}$ to be orthogonal to the gradients of the old tasks. During the learning of the $t$-th task, InfLoRA only tunes $\bm{A}_{t}$ in the new LoRA branch while freezing $\bm{B}_{t}$ and all the old LoRA branches. For further details on InfLoRA, we refer readers to the original paper [32].
@@
If the constraints in (7) are satisfied, subspaces $\{\mathcal{M}_{t,l}\}_{l=1}^{L+1}$ remain unchanged during the learning of the $t$-th task. Furthermore, for any input $\bm{x}$ from the previous t1t-1 tasks, $g_{t}(\bm{x})$ remains unchanged during the learning of the $t$-th task.
@@
For any $\bm{x}$ from previous t1t-1 tasks, we rewrite $\bm{g}_{t}(\bm{x})$ as
$g_{t}(\bm{x})=f(\bm{G}_{t,L+1}\bm{p}_{t,L}),$
$\bm{p}_{t,l}=\sigma(\bm{G}_{t,l}\bm{p}_{t,l-1}),~l\in\{1,2,...,L\},$
$\bm{p}_{t,0}=\bm{p}_{0}={\rm Pool}({\rm Token}(\bm{x})).$ (19)
Since $\bm{p}_{t,0}={\rm Pool}({\rm Token}(\bm{x}))$ is unrelated to the parameters of the new gating module $g_{t}(\cdot)$, $\bm{p}_{t,0}$ does not change with the update of $g_{t}(\cdot)$. Since $\mathcal{M}_{t,1}$ is spanned by $\bm{p}_{t,0}$, $\mathcal{M}_{t,1}$ remains unchanged during the learning of the $t$-th task.
Suppose that we have proven that $\bm{p}_{t,l-1}$ does not change with the update of the new gating module $g_{t}(\cdot)$ ($1\leq l\leq L$). Since $\mathcal{M}_{t,l}$ is spanned by $\bm{p}_{t,l-1}$, $\mathcal{M}_{t,l}$ remains unchanged during the learning of the $t$-th task. At this point, $\bm{p}_{t,l}$ can be expressed as
$\bm{p}_{t,l}=\sigma(({\rm Init}(\bm{G}_{t,l})+\Delta\bm{G}_{t,l})\bm{p}_{t,l-1})=\sigma({\rm Init}(\bm{G}_{t,l})\bm{p}_{t,l-1}).$ (20)
Here, the second equality holds since $\bm{p}_{t,l-1}\in\mathcal{M}_{t,l}$ and $\Delta\bm{G}_{t,l}\bot\mathcal{M}_{t,l}$. Therefore, $\bm{p}_{t,l}$ does not change with the update of the new gating module $g_{t}(\cdot)$ ($1\leq l\leq L$). Since $\mathcal{M}_{t,l+1}$ is spanned by $\bm{p}_{t,l}$, $\mathcal{M}_{t,l+1}$ remains unchanged during the learning of the $t$-th task.
@@
$\bm{g}_{t}(\bm{x})=f(({\rm Init}(\bm{G}_{t,L+1})+\Delta\bm{G}_{t,L+1})\bm{p}_{t,L})=f({\rm Init}(\bm{G}_{t,L+1})\bm{p}_{t,L}).$ (21)
Here, the second equality holds since $\bm{p}_{t,L}\in\mathcal{M}_{t,L+1}$ and $\Delta\bm{G}_{t,L+1}\bot\mathcal{M}_{t,L+1}$. ∎
@@ Appendix B More Details of Experimental Settings
B.1 More Details of Datasets
@@
Task639_\_multi_\_woz_\_user_\_utterance_\_generation summarization Rouge-L
Task1590_\_diplomacy_\_text_\_generation summarization Rouge-L
Task1729_\_personachat_\_generate_\_next summarization Rouge-L
Task181_\_outcome_\_extraction information extraction Rouge-L
Task748_\_glucose_\_reverse_\_cause_\_event_\_detection information extraction Rouge-L
Task1510_\_evalution_\_relation_\_extraction information extraction Rouge-L
Task002_\_quoref_\_answer_\_generation dialogue generation Rouge-L
Task073_\_commonsenseqa_\_answer_\_generation dialogue generation Rouge-L
Task591_\_sciq_\_answer_\_generation dialogue generation Rouge-L
Task511_\_reddit_\_tifu_\_long_\_text_\_summarization question answering Rouge-L
Task1290_\_xsum_\_summarization question answering Rouge-L
Task1572_\_samsum_\_summary question answering Rouge-L
Task363_\_sst2_\_polarity_\_classification sentiment analysis Accuracy
Task875_\_emotion_\_classification sentiment analysis Accuracy
Task1687_\_sentiment140_\_classification sentiment analysis Accuracy
@@
SuperNI Benchmark 1
task1572 $\rightarrow$ task363 $\rightarrow$ task1290 $\rightarrow$ task181 $\rightarrow$ task002 $\rightarrow$
task1510 $\rightarrow$ task639 $\rightarrow$ task1729 $\rightarrow$ task073 $\rightarrow$ task1590 $\rightarrow$
task748 $\rightarrow$ task511 $\rightarrow$ task591 $\rightarrow$ task1687 $\rightarrow$ task875
2
task748 $\rightarrow$ task073 $\rightarrow$ task1590 $\rightarrow$ task639 $\rightarrow$ task1572 $\rightarrow$
task1687 $\rightarrow$ task591 $\rightarrow$ task363 $\rightarrow$ task1510 $\rightarrow$ task1729 $\rightarrow$
task181 $\rightarrow$ task511 $\rightarrow$ task002 $\rightarrow$ task1290 $\rightarrow$ task875
CL Benchmark 3
MNLI $\rightarrow$ CB $\rightarrow$ WiC $\rightarrow$ COPA $\rightarrow$ QQP $\rightarrow$
BoolQA $\rightarrow$ RTE $\rightarrow$ IMDB $\rightarrow$ Yelp $\rightarrow$ Amazon $\rightarrow$
SST-2 $\rightarrow$ DBpedia $\rightarrow$ AG News $\rightarrow$ MultiRC $\rightarrow$ Yahoo
4
Yelp $\rightarrow$ Amazon $\rightarrow$ MNLI $\rightarrow$ CB $\rightarrow$ COPA $\rightarrow$
QQP $\rightarrow$ RTE $\rightarrow$ IMDB $\rightarrow$ SST-2 $\rightarrow$ DBpedia $\rightarrow$
AG News $\rightarrow$ Yahoo $\rightarrow$ MultiRC $\rightarrow$ BoolQA $\rightarrow$ WiC
B.2 More Implementation Details
Following existing CL works [40, 61, 68], all methods are implemented using instruction tuning [40]. Experiments are conducted on 5 NVIDIA RTX A6000 GPUs with AdamW [35] as the optimizer. The type of CPU is Intel(R) Xeon(R) Gold 6240R CPU @ 2.40GHz. For T5-Large and T5-XL, their relatively smaller model sizes allow experiments to be performed on a single A6000 GPU with gradient accumulation. For Llama-2-7B and Llama-2-13B, data parallelism with DeepSpeed ZeRO-2 [44] is prioritized across multiple A6000 GPUs. FlashAttention-2 [9] is employed to reduce memory usage during training, ensuring sufficient GPU memory to enable DeepSpeed ZeRO-2 whenever possible. However, if the sequence lengths of certain tasks are too long to enable DeepSpeed ZeRO-2 even with FlashAttention-2, DeepSpeed ZeRO-3 is utilized to handle these tasks.
To ensure fair comparisons, for all the methods based on LoRA, we follow existing CL methods [21, 61, 75] by integrating the LoRA architecture into the query and value components of the multi-head attention mechanism in each Transformer block. Following existing works [61, 75], for all the methods based on LoRA, the rank of a single LoRA branch is set to 4 for Order 1 and Order 2, and 8 for Order 3 and Order 4. We also vary the rank in LoRA branches and show the results in Appendix C.4.
@@ B.2 More Implementation Details
B.3 More Details about the Architecture of the Gating Module
@@
$g_{i}(\bm{x})=f(\bm{G}_{i,L+1}\bm{p}_{i,L}),$
$\bm{p}_{i,l}=\sigma(\bm{G}_{i,l}\bm{p}_{i,l-1}),~l\in\{1,2,...,L\},$
$\bm{p}_{i,0}=\bm{p}_{0}={\rm Pool}({\rm Token}(\bm{x})).$ (22)
Non-linear activation function $\sigma(\cdot)$ is set to SiLU [13]. For all experiments, unless otherwise stated, $L$ is set to 2. In other words, the gating module $g_{i}(\cdot)$ has three layers. For T5-Large and T5-XL, the parameters in the $i$-th gating module $g_{i}(\cdot)$ are $\bm{G}_{i,1}\in\mathbb{R}^{100\times d}$, $\bm{G}_{i,2}\in\mathbb{R}^{d\times 100}$ and $\bm{G}_{i,3}\in\mathbb{R}^{1\times d}$. For Llama-2-7B and Llama-2-13B, the parameters in the $i$-th gating module $g_{i}(\cdot)$ are $\bm{G}_{i,1}\in\mathbb{R}^{50\times d}$, $\bm{G}_{i,2}\in\mathbb{R}^{d\times 50}$ and $\bm{G}_{i,3}\in\mathbb{R}^{1\times d}$. Here, $d$ denotes the dimension of the embeddings. For different models, $d$ is 1024 for T5-Large and T5-XL, 4096 for Llama-2-7B, and 5120 for Llama-2-13B.
@@
B.4 Computation of Trainable Parameters
@@
B.4.1 Computation of Trainable Parameters in T5-Large
In T5-Large, the projection weights for the query and value components have shapes $\bm{W}_{q},\bm{W}_{v}\in\mathbb{R}^{1024\times 1024}$. The model consists of 24 self-attention modules in the encoder, 24 self-attention modules in the decoder, and 24 cross-attention modules in the decoder, resulting in a total of (24+24+24)2=144(24+24+24)*2=144 pre-trained weights that incorporate the LoRA architecture.
During the learning of the $t$-th new task, O-LoRA updates the parameters $\bm{A}_{t}\in\mathbb{R}^{1024\times r}$ and $\bm{B}_{t}\in\mathbb{R}^{r\times 1024}$, resulting in 1024r144+r1024144=294912r1024*r*144+r*1024*144=294912r trainable parameters. When $r=4$, the number of trainable parameters in O-LoRA is 2949124=1179648=1.18294912*4=1179648=1.18M. InfLoRA only updates the parameters $\bm{A}_{t}\in\mathbb{R}^{1024\times r}$, resulting in 1024r144=147456r1024*r*144=147456r trainable parameters. When $r=4$, the number of trainable parameters in InfLoRA is 147456r=589824=0.59147456r=589824=0.59M.
GainLoRA introduces an additional new gating module $g_{t}(\cdot)$ with parameters $\bm{G}_{t,1}\in\mathbb{R}^{100\times 1024}$, $\bm{G}_{t,2}\in\mathbb{R}^{1024\times 100}$ and $\bm{G}_{t,3}\in\mathbb{R}^{1\times 1024}$. Therefore, the number of trainable parameters in GainLoRA (O-LoRA) is 1179648+1024100+1024100+1024=1385472=1.391179648+1024*100+1024*100+1024=1385472=1.39M. The number of trainable parameters in GainLoRA (InfLoRA) is 589824+1024100+1024100+1024=795648=0.80589824+1024*100+1024*100+1024=795648=0.80M.
B.4.2 Computation of Trainable Parameters in T5-XL